OpenStudy (anonymous):

Find the soln. to the initial value problem e^x^2=xe^y with the initial condition y(0)=0

OpenStudy (anonymous):

I got the variables separated, but I have no idea how to integrate this...

OpenStudy (dumbcow):

hmm shouldn't there be a dy/dx somewhere plugging in (0,0) doesn't work for equation you posted

OpenStudy (anonymous):

oh right ! oh im srry, there's e^x^2 (y')=xe^y

OpenStudy (dumbcow):

right thats better

OpenStudy (anonymous):

:P

OpenStudy (dumbcow):

ok so you have \[\int\limits e^{-y} dy = \int\limits x e^{-x^2} dx\]

OpenStudy (anonymous):

yeah

OpenStudy (dumbcow):

left side easy .... integral of e^-y = -e^-y

OpenStudy (anonymous):

ohhhh right~ ok haha I hd it as a fraction

OpenStudy (dumbcow):

right side use substitution u = -x^2 du = -2x dx --> dx = -du/2x

OpenStudy (dumbcow):

haha yeah for integrating its usually best to get rid of fractions and use neg exponents

OpenStudy (anonymous):

ok so I've got -y on one side now, and ln[(e^-x^2)/2]....how to deal with the right side?

OpenStudy (anonymous):

OH WAIT YOU CAN SPLIT IT

OpenStudy (anonymous):

oh wait i forgot about the c

OpenStudy (anonymous):

how do I deal with the c??? TToTT

OpenStudy (anonymous):

Are you still there?

OpenStudy (dumbcow):

\[e^{-y} = \frac{1}{2} e^{-x^2} +\frac{C}{2}\] \[-y = \ln \frac{e^{-x^2} +C}{2}\]

OpenStudy (anonymous):

*magic* ok so now we just plug in a dnsolve for c?

OpenStudy (dumbcow):

right

OpenStudy (anonymous):

could you take the ln2 out, or would that be unwise and a weird hassle?

OpenStudy (dumbcow):

yes you could but i would wait til the end, its not needed to find C

OpenStudy (anonymous):

....how so?

zepdrix (zepdrix):

I feel like it might be easier to deal with the constant from this point:\[\Large\rm e^{-y} = \frac{1}{2} e^{-x^2} +c\]But maybe that's just me. Make sure you understand how to read function notation:\[\Large\rm y(0)=0 \quad means: \quad y=0 \quad when \quad x=0\]So just plug 0 in for both to solve for c.

OpenStudy (dumbcow):

x=0, y=0 e^-y = e^0 = 1 e^-x^2 = e^0 = 1 \[1 = \frac{1+C}{2}\]

OpenStudy (anonymous):

oh ok so you do need the 2 there~

OpenStudy (dumbcow):

yeah exactly ^^ i have a bad habit of trying to solve the whole thing first before finding C

OpenStudy (anonymous):

>o< haha that doesn't work out well, does it? ^^ thank you~!!! I'm gonna head in now~ you were awesome XD

OpenStudy (dumbcow):

yw :)